40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15081     Entry time: Fri Dec 6 15:22:01 2019
Author: gautam 
Type: Frogs 
Category: LSC 
Subject: DAFI system revived 

[Jordan, gautam]

We did the following:

  • Route the fiber from the control room to 1Y2.
  • Plug fiber in to FiBox at either end, turned FiBoxes ON.
  • Tested the optical connection by driving a 1Vpp 440 Hz sine wave from a function generator - Yehonathan hears it loud and clear in the control room.
  • Tested that both CH1 and CH2 work - only CH1 is connected to the speakers in the control room at the moment.
  • There is some cross-coupling between the channels - not sure if this is happening in the multi-mode fiber or in the electroncis, but I estimate the isolation to be >30dB.
  • Connected CH8 and CH9 of DAC0 in the c1lsc expansion chassis to CH1 and CH2 respectively of the FiBox in 1Y2. 
  • Restarted the c1daf model on c1lsc, came up smooth.
  • Routed the POY11 error signal through the various matrices in c1daf, and we could 👂 the Y-arm cavity 🔐 😎 
  • Channels are muted for now - I'll give this a whirl while doing the PRFPMI locking.
ELOG V3.1.3-